OpenStudy (anonymous):

Ira bought a tennis racquet that cost $112. The sales tax rate is 9 percent. What is the total amount that she paid?

OpenStudy (anonymous):

umm,i think its 10.08. 112*9=1008/100=10.08+112=122.08

OpenStudy (anonymous):

good answer...

OpenStudy (experimentx):

guess you are right

OpenStudy (experimentx):

Oo .. i thought the answerer was same as asker

OpenStudy (kymber):

Why did you bump this?

OpenStudy (anonymous):

lol i ment o close it

OpenStudy (anonymous):

Thanks!